Frank is 4 years older than Bill. Five years ago Frank was twice as old as Bill. How old is Frank now?
Answer by nyc_function(2741)

Frank's current age = x + 4

Bill's current age = x

(x + 4) - 5 = 2(x - 5)

x - 1 = 2x - 10

x - 2x = -10 + 1

-x = -9

x = -9/-1

x = 9

Frank is 9 + 4 years old or 13 years old.